OpenStudy (anonymous):

A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2, and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold.   1. Change the equation into slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all of your work.

OpenStudy (anonymous):

@pooja195

OpenStudy (ganpat):

In the slope format, Ideal equation of line is y =mx + C.. where m is the slope and C is the constant Given equation, 2x + 3y = 1200 converting it into given form , y = mx + c 3y = -2x + 1200 y = -(2/3)x + 1200/3 y = -(2/3)x + 400 there for slope is m = -(2/3) Now, to find y intercept in the above equation put x = 0 and solve for y !!